# FlashTex: Fast Relightable Mesh Texturing with LightControlNet

### v2

@@ 1 Introduction
Figure 1: We propose an efficient approach for texturing an input 3D mesh given a user-provided text prompt. Our generated texture can be relit properly in different lighting environments. The light probe shows the varied lighting environment. We suggest the readers check our video results of rotating lighting in our supplementary material.
@@ 2 Related Work
Figure 3: Our Text-to-Texture pipeline. Our method efficiently synthesizes relightable textures given an input 3D mesh and text prompt. In stage 1 (top left), we use multi-view visual prompting with our LightControlNet model to generate four visually consistent canonical views of the mesh under fixed lighting, concatenated into a reference image $I_{\text{ref}}$. In stage 2 we apply a new texture optimization procedure using $I_{\text{ref}}$ as guidance along with a multi-resolution hash-grid representation of the texture $\Gamma(\beta(\cdot))$. For each optimization iteration, we render two batches of images using $\Gamma(\beta(\cdot))$ – one using the canonical views and lighting of $I_{\text{ref}}$ to compute a reconstruction loss $\mathcal{L}_{\text{recon}}$ and the other using randomly sampled views and lighting to compute an SDS loss $\mathcal{L}_{\text{SDS}}$ based on LightControlNet.
3D texture generation. The area of 3D texture generation has evolved over time. Earlier models either directly took 3D representations as input to neural networks [4, 49, 61] or used them as templates [36, 38]. While some methods also use differentiable rendering to learn from 2D images [4, 14, 61, 38], the learned models often fail to generalize beyond the limited training categories.
Closest to our work are the recent works that use pre-trained 2D diffusion models and treat texture generation as a byproduct of text-to-3D generation. Examples include Latent-Paint [29], which uses Score Distillation Sampling in latent space, Text2tex [8], which leverages depth-based 2D ControlNet, and TEXTure [43], which exploits both previous methods. Nonetheless, similar to recent text-to-3D models, such methods produce textures with entangled lighting effects and suffer from slow generation. On the other hand, TANGO [10], generates material textures using a Spherical-Gaussian-based differentiable renderer, but struggles with complex texture generation. A concurrent work, Paint3D [62], aims to generate lighting-less textures, yet it cannot produce material-based textures like ours.
@@ 3 Preliminaries
Score Distillation Sampling (SDS). DreamFusion [39] optimizes a 3D scene representation conditioned on text prompts using a pre-trained 2D text-to-image diffusion model. The scene is represented as a NeRF  [30, 1] parametrization $\theta$. A differentiable renderer $\mathcal{R}$ applied to $\theta$ with a randomly sampled camera view $C$ then generates a 2D image $x=\mathcal{R}(\theta,C)$. A small amount of noise $\epsilon\sim\mathbb{N}(0,1)$ is then added to $x$ to obtain a noisy image $x_{t}$. DreamFusion leverages a diffusion model $\phi$ (Imagen [45]) to provide a score function $\hat{\epsilon}_{\phi}(x_{t};y,t)$, which predicts the sampled noise $\epsilon$ given the noisy image $x_{t}$, text prompt $y$, and noise level $t$. This score function can update the scene parameters $\theta$, using the gradient calculated by Score Distillation Sampling (SDS):
@@
where $w(t)$ is a weighting function. During each iteration, to calculate the SDS loss, we randomly choose a camera view $C$, render the NeRF $\theta$ to form an image $x$, add noise $\epsilon$ to it, and predict the noise using the diffusion model $\phi$. We run the optimization for 5,000 to 10,000 iterations.
In our work, we introduce an illumination-aware SDS loss to optimize surface texture on a 3D mesh to suppress inconsistency artifacts and simultaneously separate lighting from the surface reflectance.
@@ 4 Method
Our text-to-texture pipeline operates in two main stages to generate a relightable texture for an input 3D mesh with a corresponding text prompt (Figure 3). In Stage 1, we use a multi-view visual prompting approach to obtain visually consistent views of the object from a small set of viewpoints, using a 2D ControlNet. Simply backprojecting these sparse views onto the 3D mesh could produce patches of high-quality texture, but would also generate visible seams and other visual artifacts where the views do not fully match. The resulting texture would also have lighting baked-in, making it difficult to relight the textured mesh in a new lighting environment. Therefore, in Stage 2, we apply a texture optimization that uses a ControlNet in combination with Score Distillation Sampling (SDS) [39] to mitigate such artifacts and separate lighting from the surface material properties/reflectance. In both stages, we introduce a new illumination-aware ControlNet that allows us to specify the desired lighting as a conditioning image for an underlying text-to-image diffusion model. We call this model LightControlNet and describe how it works in Section 4.1. We then detail each stage in Section 4.2 and Section 4.3, respectively.
Figure 4: (a) LightControlNet requires a conditioning image that specifies desired lighting $L$ for a view $C$ of a 3D mesh. To form the conditioning image, we first render the mesh with the desired $L$ and $C$ using three different materials: (1) non-metal, not smooth, (2) half-metal, half-smooth, and (3) pure metal, smooth, and then combine the renderings into a single three-channel image. (b) LightControlNet is a diffusion model that is conditional on such light conditioning images as well as text prompts.
@@ 4.1 LightControlNet
LightControlNet adapts the ControlNet architecture to enable control over the lighting in the generated image. More specifically, we create a conditioning image for a 3D mesh by rendering it using three pre-defined materials and under known lighting conditions (Figure 4). These renderings encapsulate information about the desired shape and lighting for the object, and we stack them into a three-channel conditioning image. We have found that setting the pre-defined materials to (1) non-metal, not smooth; (2) half-metal, half-smooth; and (3) pure metal, extremely smooth, respectively, works well in practice. The specific material parameters are in the appendix.
@@
Distilling the encoder. We improve the efficiency of LightControlNet by distilling the image encoder in Stable Diffusion [44], the base diffusion model in the ControlNet architecture. The original Stable Diffusion image encoder consumes almost 50% of the forward and backward time of SDS calculation using the latent diffusion model, primarily in downsampling the input image. Metzer et al. [29] have found the image decoder from latent space to image space can be closely approximated by per-pixel matrix multiplication. Inspired by this, we distill the encoder by removing its attention modules and training it on the COCO dataset [26] to match the original output. This distilled encoder runs 5x faster than the original one, resulting in an approximately 2x acceleration of our text-to-texture pipeline without compromising output quality. An ablation study of our distilled encoder is detailed in Table 3, with additional implementation specifics in the appendix.
@@ 4.2 Stage 1: Multi-view Visual Prompting
To mitigate this multi-view inconsistency issue, we take a multi-view visual prompting approach. We concatenate the conditioning images for the four canonical views into a single 2 $\times$ 2 grid and treat it as a single conditioning image. We observe that applying LightControlNet to all four views simultaneously, using this combined multi-view conditioning image, results in a far more consistent appearance and style across the views, compared to independent prompting (Figure 5). We suspect this property arises from the presence of similar training data samples – grid-organized sets depicting the same object – in Stable Diffusion’s training set, which is also observed in concurrent works [59, 64]. Formally, we generate the conditioning image $I_{\text{cond}}(L^{*},C^{*})$ under a fixed canonical lighting condition $L^{*}$ using four canonical viewpoints $C^{*}$. We then apply our LightControlNet with text prompt $y$ to generate the corresponding reference image $I_{\text{ref}}$:
@@ 4.3 Stage 2: Texture Optimization
In Stage 2, we could directly backproject the four reference views output in Stage 1 onto the 3D mesh using the camera matrix $C$ associated with each view. While the resulting texture would contain some high-quality regions, it would also suffer from two problems (1) It would contain seams and visual artifacts due to remaining inconsistencies between overlapping views, occlusions in the views that leave parts of the mesh untextured, and loss of detail when applying the backprojection transformation and resampling the views. (2) In addition, as lighting is baked into the LightControlNet’s RGB images, it would also be baked into the backprojected texture, making it difficult to relight the mesh.
To address both of these issues, we employ texture optimization using SDS loss. Specifically, we use a multi-resolution hash-grid [32] as our 3D scene representation, instead of NeRF as in the original DreamFusion formulation [39]. Given a 3D point $p\in\mathbb{R}^{3}$ on the mesh, our hash-grid produces a 32-dimensional multi-resolution feature. This feature is then fed to a 2-layer MLP $\Gamma$ to obtain the texture material parameters for this point. Similar to Fantasia3D [9], these material parameters consist of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ and the base color $k_{c}\in\mathbb{R}^{3}$. Formally,
@@
Scheduling the optimization. We warm up the optimization by rendering the four canonical views and applying $\mathcal{L}_{\text{recon}}$ for 50 iterations. We then add in iterations using the $\mathcal{L}_{\text{SDS}}$ loss and optimize over randomly chosen camera views and randomly selected lighting from a pre-defined set of environmental lighting maps. Specifically we alternate iterations between using $\mathcal{L}_{\text{SDS}}$ and $\mathcal{L}_{\text{recon}}$. In addition, for a quarter of the SDS iterations, we use the canonical views rather than randomly selecting the views. This ensures that the resulting texture does not overfit to the reference images corresponding to the canonical views. The warm-up iterations capture the large-scale structure of our texture and allow us to use relatively small noise levels ($t\leq 0.1$) in the SDS optimization. We sample the noise following a linearly decreasing schedule [17] with $t_{\text{max}}=0.1$ and $t_{\text{min}}=0.02$. We also adjust the conditioning strength $s$ of our LightControlNet in the SDS loss linearly from $1$ to 00 over these iterations so that LightControlNet is only lightly applied by the end of the optimization. We also experimented with a recent variant Variational Score Distillation [57], but did not observe notable improvement. We have experimentally found that we obtain high-quality textures after 400 total iterations of this optimization and this is far fewer iterations than other SDS-based texture generation techniques such as Fantasia3D [9] which requires 5000 iterations. More details are in the appendix.
@@ 5 Experiments
Dataset. As illustrated in Figure 3, we employ Objaverse [11] to render paired data to train our LightControlNet. Objaverse consists of approximately 800k objects, of which we use the names and tags as their text descriptions. We filter out objects with low CLIP similarity [40] to their text descriptions and select around 40k as our training set. Each object is rendered from 12 views using randomly sampled cameras and lighting from a specific set of environmental lighting maps. To evaluate baselines and our method, we hold out 70 random meshes from Objaverse [11] as the test set. We additionally gather 22 mesh assets from 3D online games with 5 prompts each to assess our method, demonstrating its ability to generalize beyond Objaverse. Further details are in the appendix.
@@
Figure 7: Qualitative analysis. (a) We compare our method with Fantasia3D [9] that also attempts to generate Physically Based Rendering (PBR) texture. However, unlike ours, their results often exhibit baked-in lighting, leading to artifacts when put into varied lighting environments. (b) We also compare our method with other baselines that can only generate non-relightable (RGB) texture. For non-relightable texture generation, we can replace our LightControlNet with depth ControlNet and generate RGB textures with a shorter runtime. More details are in Table 1.
@@
Quantative evaluation. In Table 1, we compare our method with the baselines on the Objaverse [11] test set. For each method, we generate 16 views and evaluate Frechet Inception Distance (FID) [16, 37] and Kernel Inception Distance (KID) [3] compared with ground-truth rendered views. Two variations of our method are assessed. Both variants use our proposed two-stage pipeline, and the first employs a standard depth-guided ControlNet, while the second uses our proposed LightControlNet. Our method significantly outperforms the baselines in both quality and runtime.
Qualitative analysis. As shown in Figure 6, our method can generate highly-detailed textures that can be rendered properly with the environment lighting across a wide variety of meshes. We also visually compare our method and the baselines in Figure 7. Our method produces textures with higher visual fidelity than the baselines for both the relightable and non-relightable variants. In particular, when compared with Fantasia3D [9], a recent work that also aims to generate material-based texture, our results not only have superior visual quality, but also disentangle the lighting more successfully.
User study. To further evaluate the texture quality quantitatively, we conduct a user study comparing our results with each of the baselines on the Objaverse test set in Table 2. We ask 30 participants to evaluate (1) the realism of the results, (2) the consistency of the generated texture with the input text, and (3) the plausibility of the results when placed under varying lighting conditions. Each result is presented in the form of 360-degree rotation to display full texture details. The reference lighting is provided alongside when participants evaluate (3). Across all three aspects, participants consistently prefer our method over baselines.
Table 3: Ablation study on algorithmic components. We analyze the role of our distilled encoder (1st row) and multi-view visual prompting (2nd row). Replacing the distilled encoder with the original VQ-VAE encoder doubles the running time without a noticeable improvement. When removing the multi-view visual prompting for initial generation, the system requires 2,000 iterations (5x compared to our 400 iterations) to produce reasonable results, which produces slightly worse texture quality.
@@
Table 5: Ablation study on the number of canonical views. We analyze the role of our canonical view selection in Section 4.2. Relying on only the left and right views provides insufficient supervision. Interestingly, adding top and bottom views leads to worse overall quality. This is likely due to the limitation of pre-trained 2D diffusion models in synthesizing top and bottom views well for a variety of objects. Furthermore, given the fixed resolution of the multi-view image, stacking more views would result in a lower resolution for each view, leading to a worse initialization for Stage 2.
@@
Ablation study. We perform a thorough ablation analysis on different aspects of our method as seen in Table 3. When substituting our distilled encoder with the original VQ-VAE encoder, the performance is twice as slow, but the quality of results is not noticeably superior. On the other hand, without the multi-view visual prompting for the initial generation, the system requires 2000 iterations (a 5x slowdown compared to our 400 iterations) to produce reasonable results, while still leading to slightly worse texture quality.
In Section 4.1, we render a conditioning image using 3 pre-defined materials to encompass a broad range of feasible materials: (1) non-metal, not smooth (diffuse effect); (2) half-metal, half-smooth (mixed effect); (3) pure metal, smooth (specular effect). These material bases cover a large range of feasible materials. Table 4 shows omitting any one of these bases degrades quality.
As shown in Table 5, we also evaluate our selection of four canonical views in Section 4.2. Relying on only the left and right views provides insufficient supervision. Interestingly, incorporating top and bottom views degrades the performance. We hypothesize that the degradation is likely due to the limitation of 2D diffusion model backbones in generating top and bottom views reliably. Furthermore, stacking more views within a single image results in a decreased resolution for each view, given the fixed resolution of the multi-view image.

### v3

@@ 1 Introduction
Figure 1: We propose an efficient approach for texturing an input 3D mesh given a user-provided text prompt. Our generated texture can be relit properly in different lighting environments. The light probe shows the varied lighting environment. We suggest the readers check our video results of rotating lighting at our website.
@@ 2 Related Work
Figure 3: Text-to-Texture pipeline. Our method efficiently synthesizes relightable textures given a 3D mesh and text prompt. In stage 1 (top left), we use multi-view visual prompting with our LightControlNet to generate four visually consistent canonical views of the mesh under fixed lighting, concatenated into a reference image $I_{\text{ref}}$. In stage 2, we apply a new texture optimization procedure using $I_{\text{ref}}$ as guidance along with a multi-resolution hash-grid representation of the texture $\Gamma(\beta(\cdot))$. For each iteration, we render two batches of images using $\Gamma(\beta(\cdot))$ – one using the canonical views and lighting of $I_{\text{ref}}$ to compute a reconstruction loss $\mathcal{L}_{\text{recon}}$ and the other using randomly sampled views and lighting to compute an SDS loss $\mathcal{L}_{\text{SDS}}$ based on LightControlNet.
3D texture generation. The area of 3D texture generation has evolved over time. Earlier models either directly took 3D representations as input to neural networks [4, 49, 61] or used them as templates [36, 38]. While some methods also use differentiable rendering to learn from 2D images [4, 14, 61, 38], the models often fail to generalize beyond the limited training categories. Closest to our work are the recent works that use pre-trained 2D diffusion models and treat texture generation as a byproduct of text-to-3D generation. Examples include Latent-Paint [29], which uses Score Distillation Sampling in latent space, Text2tex [8], which leverages depth-based 2D ControlNet, and TEXTure [43], which exploits both previous methods. Nonetheless, similar to recent text-to-3D models, such methods produce textures with entangled lighting effects and suffer from slow generation. On the other hand, TANGO [10], generates material textures using a Spherical-Gaussian-based differentiable renderer, but struggles with complex texture generation. A concurrent work, Paint3D [62], aims to generate lighting-less textures, yet it cannot produce material-based textures like ours.
@@ 3 Preliminaries
Score Distillation Sampling (SDS). DreamFusion [39] optimizes a 3D NeRF representation $\theta$ [30, 1] conditioned on text prompts using a pre-trained 2D text-to-image diffusion model. A differentiable renderer $\mathcal{R}$ applied to $\theta$ with a randomly sampled camera view $C$ then generates a 2D image $x=\mathcal{R}(\theta,C)$. A small amount of noise $\epsilon\sim\mathbb{N}(0,1)$ is then added to $x$ to obtain a noisy image $x_{t}$. DreamFusion leverages a diffusion model $\phi$ (Imagen [45]) to provide a score function $\hat{\epsilon}_{\phi}(x_{t};y,t)$, which predicts the sampled noise $\epsilon$ given the noisy image $x_{t}$, text prompt $y$, and noise level $t$. This score function can update the scene parameters $\theta$, using the gradient calculated by SDS:
@@
where $w(t)$ is a weighting function. During each iteration, to calculate the SDS loss, we randomly choose a camera view $C$, render the NeRF $\theta$ to form an image $x$, add noise $\epsilon$ to it, and predict the noise using the diffusion model $\phi$. DreamFusion optimzes for 5,000 to 10,000 iterations. In our work, we introduce an illumination-aware SDS loss to optimize surface texture on a mesh to suppress inconsistency artifacts and simultaneously separate lighting from the material.
@@ 4 Method
Our text-to-texture pipeline operates in two main stages to generate a relightable texture for an input 3D mesh with a text prompt (Figure 3). In Stage 1, we use multi-view visual prompting to obtain visually consistent views of the object from a small set of viewpoints, using a 2D ControlNet. Simply backprojecting these sparse views onto the 3D mesh could produce patches of high-quality texture but would also generate visible seams and other visual artifacts where the views do not fully match. The resulting texture would also have lighting baked-in, making it difficult to relight the textured mesh in a new lighting environment. Therefore, in Stage 2, we apply a texture optimization that uses a ControlNet in combination with Score Distillation Sampling (SDS) [39] to mitigate such artifacts and separate lighting from the surface material properties. In both stages, we introduce a new illumination-aware ControlNet that allows us to specify the desired lighting as a conditioning image for an underlying text-to-image diffusion model. We call this model LightControlNet and describe how it works in Section 4.1. We then detail each stage in Section 4.2 and Section 4.3, respectively.
Figure 4: (a) LightControlNet requires a conditioning image that specifies desired lighting $L$ for a view $C$ of a 3D mesh. To form the conditioning image, we render the mesh with the desired $L$ and $C$ using three different materials: (1) non-metal, not smooth, (2) half-metal, half-smooth, and (3) pure metal, smooth, and then combine the renderings into a single three-channel image. (b) LightControlNet is a diffusion model conditioned on such light-conditioning images and text prompts.
@@ 4.1 LightControlNet
LightControlNet adapts the ControlNet architecture to enable control over the lighting in the generated image. Specifically, we create a conditioning image for a 3D mesh by rendering it using three pre-defined materials and under known lighting conditions (Figure 4). These renderings encapsulate information about the desired shape and lighting for the object, and we stack them into a three-channel conditioning image. We have found that setting the pre-defined materials to (1) non-metal, not smooth; (2) half-metal, half-smooth; and (3) pure metal, extremely smooth, respectively, works well in practice.
@@
Distilling the encoder. We improve the efficiency of SDS by distilling the image encoder in Stable Diffusion (SD) [44], the base diffusion model in the ControlNet architecture. The original SD encoder consumes almost 50% of the forward and backward time of SDS calculation, primarily in downsampling the input image. Metzer et al. [29] have found the image decoder can be closely approximated by per-pixel matrix multiplication. Inspired by this, we distill the encoder by removing its attention modules and training it on the COCO dataset [26] to match the original output. This distilled encoder runs 5x faster than the original one, resulting in an approximately 2x acceleration of our text-to-texture pipeline without compromising output quality (Table 3).
@@ 4.2 Stage 1: Multi-view Visual Prompting
We use a multi-view visual prompting approach to mitigate this multi-view inconsistency issue. We concatenate the conditioning images for the four canonical views into a single 2 $\times$ 2 grid and treat it as a single conditioning image. We observe that applying LightControlNet to all four views simultaneously, using this combined multi-view conditioning image, results in a far more consistent appearance and style across the views, compared to independent prompting (Figure 5). We suspect this property arises from the presence of similar training data samples – grid-organized sets depicting the same object – in Stable Diffusion’s training set, which is also observed in concurrent works [59, 64]. Formally, we generate the conditioning image $I_{\text{cond}}(L^{*},C^{*})$ under a fixed canonical lighting condition $L^{*}$ using four canonical viewpoints $C^{*}$. We then apply our LightControlNet with text prompt $y$ to generate the corresponding reference image $I_{\text{ref}}$:
@@ 4.3 Stage 2: Texture Optimization
In Stage 2, we could directly backproject the four reference views output in Stage 1 onto the 3D mesh using the camera matrix $C$ associated with each view. While the resulting texture would contain some high-quality regions, it would also suffer from two problems: (1) It would contain seams and visual artifacts due to remaining inconsistencies between overlapping views, occlusions in the views that leave parts of the mesh untextured, and loss of detail when applying the backprojection transformation and resampling the views. (2) as lighting is baked into the LightControlNet’s RGB images, it would also be baked into the backprojected texture, making it difficult to relight the mesh.
To address both issues, we employ texture optimization using SDS loss. Specifically, we use a multi-resolution hash-grid [32] as our 3D texture representation. Given a 3D point $p\in\mathbb{R}^{3}$ on the mesh, our hash-grid produces a 32-dim multi-resolution feature, which is then fed to a 2-layer MLP $\Gamma$ to obtain the texture material parameters for this point. Similar to Fantasia3D [9], these material parameters consist of metallicness $k_{m}\in\mathbb{R}$, roughness $k_{r}\in\mathbb{R}$, a bump vector $k_{n}\in\mathbb{R}^{3}$ and the base color $k_{c}\in\mathbb{R}^{3}$. Formally,
@@
Scheduling the optimization. We warm up the optimization by rendering the four canonical views and applying $\mathcal{L}_{\text{recon}}$ for 50 iterations. We then add in iterations using $\mathcal{L}_{\text{SDS}}$ and optimize over randomly chosen camera views and randomly selected lighting from a pre-defined set of environmental lighting maps. Specifically we alternate iterations between using $\mathcal{L}_{\text{SDS}}$ and $\mathcal{L}_{\text{recon}}$. In addition, for a quarter of the SDS iterations, we use the canonical views rather than randomly selecting the views. This ensures that the resulting texture does not overfit to the reference images corresponding to the canonical views. The warm-up iterations capture the large-scale structure of our texture and allow us to use relatively small noise levels ($t\leq 0.1$) in the SDS optimization. We sample the noise following a linearly decreasing schedule [17] with $t_{\text{max}}=0.1$ and $t_{\text{min}}=0.02$. We also adjust the conditioning strength $s$ of our LightControlNet in $\mathcal{L}_{\text{SDS}}$ linearly from $1$ to 00 over these iterations so that LightControlNet is only lightly applied by the end of the optimization. We also experimented with a recent variant Variational Score Distillation [57], but did not observe notable improvement. We have experimentally found that we obtain high-quality textures after 400 total iterations of this optimization and this is far fewer iterations than other SDS-based texture generation techniques such as Fantasia3D [9] which requires 5000 iterations.
@@ 5 Experiments
Dataset. As illustrated in Figure 3, we employ Objaverse [11] to render paired data to train our LightControlNet. Objaverse consists of approximately 800k objects, of which we use the names and tags as their text descriptions. We filter out objects with low CLIP similarity [40] to their text descriptions and select around 40k as our training set. To evaluate baselines and our method, we hold out 70 random meshes from Objaverse [11] as the test set. We additionally gather 22 mesh assets from 3D online games with 5 prompts each to assess our method.
@@
Figure 7: Qualitative analysis. (a) We compare with Fantasia3D [9] that also attempts to generate Physically Based Rendering (PBR) texture. However, their results often exhibit baked-in lighting, leading to artifacts when put under varied lighting. (b) We also compare with other baselines that only generate non-relightable (RGB) texture. For non-relightable texture generation, we can replace our LightControlNet with depth ControlNet and generate textures with a shorter runtime. More details are in Table 1.
@@
Quantative evaluation. In Table 1, we compare our method with the baselines on the Objaverse [11] test set. For each method, we generate 16 views and evaluate Frechet Inception Distance (FID) [16, 37] and Kernel Inception Distance (KID) [3] compared with ground-truth renderings. Two variations of our method are assessed. Both variants use ourr two-stage pipeline, and the first employs a standard depth ControlNet, while the second uses our proposed LightControlNet. Our method outperforms the baselines in both quality and runtime.
Qualitative analysis. In Figure 6, our method can generate highly detailed textures that can be rendered properly with the environment lighting across various meshes. We also visually compare our method and the baselines in Figure 7. Our method produces textures with higher visual fidelity for both the relightable and non-relightable variants. In particular, when compared with Fantasia3D [9], a recent work that also aims to generate material-based texture, our results not only have superior visual quality, but also disentangle the lighting more successfully.
User study. To further evaluate the texture quality quantitatively, we conduct a user study comparing our results with each of the baselines on the Objaverse test set in Table 2. We asked 30 participants to evaluate (1) the realism of the results, (2) the consistency of the generated texture with the input text, and (3) the plausibility of the results when placed under varying lighting conditions. Each result is presented in the form of 360-degree rotation to display full texture details. The reference lighting is provided alongside when participants evaluate (3). Across all three aspects, participants consistently prefer our method.
Table 3: Ablation study on algorithmic components. We analyze the role of our distilled encoder (1st row) and multi-view visual prompting (2nd row). Replacing the distilled encoder with the original encoder doubles the running time without a noticeable improvement. When removing the multi-view visual prompting for initial generation, the system requires 2,000 iterations (5x compared to our 400 iterations) to produce reasonable results, which produces slightly worse texture quality.
@@
Table 5: Ablation study on canonical view selection in Section 4.2. Using only front and back views provides insufficient supervision while adding top and bottom views worsens quality. This likely stems from pre-trained 2D diffusion models struggling with top and bottom views. Additionally, stacking more views reduces each view’s resolution, leading to poorer initialization for Stage 2.
@@
Ablation study. We perform an ablation analysis on different aspects of our method in Table 3. When replacing our distilled encoder with the original SD encoder, performance is twice as slow without noticeably superior quality. On the other hand, without the multi-view visual prompting for the initial generation, the system requires 2000 iterations (a 5x slowdown compared to our 400 iterations) to produce reasonable results while still leading to slightly worse texture quality. In Section 4.1, we render a conditioning image using three pre-defined materials to encompass a broad range of feasible effects. Table 4 shows omitting any one of these bases degrades quality. Table 5 evaluates our selection of four canonical views in Section 4.2. Relying on only the front and back views provides insufficient supervision. Interestingly, incorporating top and bottom views degrades the performance. We hypothesize that this is likely due to the limitation of 2D diffusion model backbones in reliably generating top and bottom views. Furthermore, stacking more views within a single image results in a decreased resolution for each view, given the fixed resolution of the multi-view image.